# Choi-Defined Resource Theories

### v2

@@ Abstract
Many resource theories share an interesting property: An operation is free if and only if its renormalized Choi matrix is a free state. In this letter, we refer to resource theories exhibiting this property as Choi-defined resource theories. We demonstrate how and under what conditions one can construct a Choi-defined resource theory, and prove that when such a construction is possible, the free operations are all and only the completely resource non-generating operations. Moreover, we define resource measures, a complete family of monotones, and conversion distances in such resource theories.
@@
The recent booming interest in quantum computing, communication, and information originates from the understanding that quantumness allows us to outperform the classical world, which means that quantum objects are resources. The new technological developments, which are now becoming commercially available, demand a rigorous and thorough understanding of quantum resources, in order to quantify them precisely, and optimize their use. This is accomplished with the framework of resource theories [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. Resource theories are based on the simple observation that agents in a laboratory can only perform a subset of all possible deterministic quantum operations due to the various constraints on the agents and the systems involved. For instance, when two agents, Alice and Bob, are physically separated, they may not have a reliable means to exchange quantum systems, limiting them to local operations and classical communication (LOCC) [5, 6, 7]. Such allowed operations are called free.
@@
Remarkably, one notices that in each of these instances of resource theories a channel is considered free if and only if its renormalized Choi matrix is a free state. We call resource theories that exhibit this property Choi-defined resource theories. This is a highly valuable property: Every problem associated with free quantum channels can be converted into a problem involving free states, which, in turn, can be conveniently tackled with tools developed in linear algebra, matrix analysis, and convex optimization.
Given the significance of Choi-defined resource theories and their computational convenience one wonders when it is possible to construct such resource theories from a set of free states. In this letter, we answer this question by providing necessary and sufficient conditions on the set of free states. Moreover, we show that when such a construction is possible, the resulting set of free operations coincides with the set of CRNG operations associated with those free states. These results provide a constructive definition of CRNG operations and generalize what has been already observed in the resource theories of magic, imaginarity, NPT, SEP, and non-negativity of quantum amplitudes. Next, we introduce resource measures, a complete family of monotones, and conversion distances in CDRTs, which can all be computed with conic linear programs (CLP) [48, 49] whenever the set of free states is convex and closed, or even with semidefinite programs (SDP) [50]. Lastly, we present the resource theory of asymmetry [51], for which the correspondence between the free states and the Choi matrices of the free channels was not already known, and we also characterize which resource theories of asymmetries are CDRTs.
@@
    parallel composition of free channels is free
@@
The renormalized Choi state is not the maximally entangled state in the resource theories of entanglement, SEP entanglement, or NPT entanglement. Indeed, in these resource theories, each system is a pair of pysical systems [2], i.e., $S\coloneqq AB$, and $S^{\prime}\coloneqq A^{\prime}B^{\prime}$ and the spatial separation is between systems $AA^{\prime}$ and $BB^{\prime}$. Let {|$\Set{\Ket{a}_{A}}$ and {|$\Set{\Ket{b}_{B}}$ be orthogonal basis for $A$ and $B$, respectively, then
$\frac{1}{d_{S}}\Phi_{SS^{\prime}}=\frac{1}{d_{A}d_{B}}\sum_{a,b}|ab\rangle\!\langle ab|_{AB}\otimes|ab\rangle\!\langle ab|_{A^{\prime}B^{\prime}}.$ (4)
@@
$\frac{1}{d_{A}d_{B}}\sum_{a}|aa\rangle\!\langle aa|_{AA^{\prime}}\otimes\sum_{b}|bb\rangle\!\langle bb|_{BB^{\prime}}.$ (5)
@@
§ V.  Quantifying resources in CDRTs
@@
when $\operatorname{supp}\rho_{A}\subseteq\operatorname{supp}\sigma_{A}$ and +$\infty$+ ∞ otherwise. The extension of the max-relative entropy to quantum channels [58, 59] satisfies [60]
$D_{\max}(\mathcal{M}_{A\to B}\|\mathcal{N}_{A\to B})=D_{\max}(\mu_{BA}\|\nu_{BA}),$ (8)
where, $\mu_{BA}$ and $\nu_{BA}$ are the renormalized Choi matrices of $\mathcal{M}_{A\to B}$ and $\mathcal{N}_{A\to B}$, respectively. In a CDRT, this property allows us to define resource measures for states [1] and channels [61] that depend uniquely on the set of free states:
$D_{\max}^{\mathfrak{F}}(\rho_{A})=\inf_{\omega_{A}\in\mathfrak{F}(A)}D_{\max}(\rho_{A}\|\omega_{A}),$ (9)
$D_{\max}^{\mathfrak{F}}(\mathcal{M}_{A\to B})=\inf_{\begin{subarray}{c}\omega_{BA}\in\mathfrak{F}(BA),\\ d_{A}\operatorname{Tr}_{B}\omega_{BA}=\operatorname{\mathds{1}}_{A}\end{subarray}}D_{\max}(\mu_{BA}\|\omega_{BA}),$
where $\mathfrak{F}(A)$ denotes the set of free states of system $A$. Both these resource measures are finite in every CDRTs because for every $\rho_{A}$, $\operatorname{supp}\rho_{A}\subseteq\operatorname{supp}\operatorname{\mathds{1}}_{A}$ and $\frac{1}{d_{A}}\operatorname{\mathds{1}}_{A}$ is free. The same is not true for all resource theories. Moreover, these resource measures can be computed with a CLP whenever the set of free states is convex and closed, and with a SDP whenever the membership problem for the set of free states is an SDP (see the SM).
As a direct consequence of the results in Ref. [60], the resource measures in Eq. (9) satisfy
@@
for all states $\rho_{A}$ and channels $\mathcal{M}_{A\to B}$. For the reasons mentioned above, this inequality, which is valid for all resource theories, is particularly meaningful in the context of CDRTs. When we interpret $D^{\mathfrak{F}}$ as the value of a resource, it implies that value of the output of a quantum operation applied to a quantum state is always less than the sum of the values of the operation and the input state. Equivalently, for every $\rho_{A}$, $\sigma_{B}$, and $\mathcal{M}_{A\to B}$ such that $\sigma_{B}=\mathcal{M}_{A\to B}(\rho)$, the cost of preparing $\sigma_{B}$ is always less or equal than the cost of preparing $\rho_{A}$ and then converting it to $\sigma_{B}$ with the channel $\mathcal{M}_{A\to B}$. An inequality similar to the one in Eq. (10) was derived in Ref. [62]. The divergences that appear in that inequality are the relative entropy and the regularized channel relative entropy. Instead, in our case, only the max-relative entropy is used, and no regularization is needed.
§ VI.  Resource conversion in CDRTs
@@
Being a complete family of monotones, it satisfies $f_{\tau_{B}}(\rho_{A})\geq f_{\tau_{B}}(\sigma_{B})$ for all density matrices $\tau_{B}$ if and only if $\rho_{A}$ can be converted to $\sigma_{B}$ with a free operation. The advantage of this approach is that the optimization is over states and not channels. In particular, the quantity above can be computed with a CLP (see the SM).
@@
$=\inf_{\begin{subarray}{c}\mu_{BA}\in\mathfrak{F}(BA),\\ d_{A}\operatorname{Tr}_{B}\mu_{BA}=\operatorname{\mathds{1}}_{A}\end{subarray}}D(\sigma_{B},d_{A}\mu_{B\mathord{:}A^{\prime}}*\rho_{A}),$
where $D(\cdot,\cdot)$ is a function on quantum states that satisfies the data processing inequality [1]. Once again, the minimization is only over states and this can simplify the computation of the conversion distances. In particular, many metrics, such as the trace distance or the purified distance, can be computed with SDPs [50]. Consequently, the conversion distances based on such metrics are SDPs too whenever the membership problem for the set of free states can be expressed as an SDP, e.g., in the case of the resource theories of imaginarity and PPT.
§ VII.  Examples of CDRTs
As mentioned in the introduction, known results about the resource theories of SEP, PPT, magic, imaginarity, and non-negative quantum amplitudes imply that these resource theories are CDRTs and we show in the SM that their sets of free states satisfy the conditions of Theorem 3.
Of more interest are those resource theories where the identification of free operations with CD operation has not been already established. For example, to the best of our knowledge, there is no result yet in this direction for the resource theory of asymmetry. In this section, we derive under what conditions resource theories of asymmetry [51, 14] are CDRTs.
In the resource theory of asymmetry, two parties, Alice and Adam have access to the same quantum systems but they do not share a reference frame. The relation between reference frames is described by an element $g$ of a compact group $G$ identified by the symmetries of the problem. If Alice and Adam do not know how their reference frames are related, then Alice can prepare only states for Adam that are $G$-invariant, that is:
$\rho_{A}\,\text{is free}\,\Leftrightarrow\mathcal{U}_{g}^{A}(\rho_{A})=\rho_{A},\,\forall g\in G,$ (13)
where $\mathcal{U}_{g}^{A}(\rho_{A})\coloneq U^{A}_{g}\rho_{A}(U^{A}_{g})^{\dagger}$, and $U_{g}$ is the unitary representation of $g$ on $A$. Similarly, if Adam asks Alice to perform a deterministic quantum operation, then she can only perform $G$-covariant channels:
$\mathcal{M}_{A\to B}\,\text{is free}\,\Leftrightarrow\mathcal{U}^{B}_{g}\circ\mathcal{M}_{A\to B}\circ(\mathcal{U}_{g}^{A})^{\dagger}=\mathcal{M}_{A\to B}$ (14)
for all $g\in G$.
Every different finite or compact Lie group $G$ generates a different resource theory, and the following proposition characterizes the ones that are CDRTs.
Proposition 6.
A resource theory of asymmetry is a CDRT if and only if the renormalized Choi state is free.
We point out that this is equivalent to asking that all the unitary representations are real (see the SM for details). An example of such resource theory is the parity resource theory generated by the group $\mathbb{Z}_{2}$ [51].
§ VIII.  Conclusions
In this work, we identified a key property of the resource theories of magic, imaginarity, SEP, NPT, and non-negativity of quantum amplitudes, which is that a channel is considered free if and only if its suitably renormalized Choi matrix is a free state. We presented a new way to construct a resource theory from the set of free states based on the Choi isomorphism and we derived which conditions on the free states are necessary and sufficient for constructing a mathematically consistent Choi-defined resource theory. Our interest is driven by the significant advantages that such a property guarantees. These advantages stem from the fact that every question in Choi-defined resource theories can be naturally framed as a question involving free states alone, which often have a well-understood structure. Indeed, we introduce several resource quantifiers which can all be computed with CLPs if the set of free states is convex and closed. Moreover, we showed that in any CDRT free operations are all CRNG operations. This provides a new constructive way to define CRNG operations when the set of free states satisfy the conditions of Theorem 3.
@@ Acknowledgements.
The authors are grateful to Matt Wilson for the several in-depth discussions on the subject of this letter and to Gilad Gour and Nathaniel Johnston for the interesting conversations about this research. E. Z. acknowledges support from the Alberta Graduate Excellence Scholarship (AGES) and from the Eyes High International Doctoral Scholarship. C. M. S. acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) through the Discovery Grant “The power of quantum resources” RGPIN-2022-03025 and the Discovery Launch Supplement DGECR-2022-00119.

### v3

@@ Abstract
Many resource theories share an interesting property: An operation is free if and only if its renormalized Choi matrix is a free state. In this letter, we refer to resource theories exhibiting this property as Choi-defined resource theories. We demonstrate how and under what conditions one can construct a Choi-defined resource theory, and we prove that when such a construction is possible, the free operations are all and only the completely resource non-generating operations. Moreover, we examine resource measures, a complete family of monotones, and conversion distances in such resource theories.
@@
The recent booming interest in quantum computing, communication, and information originates from the understanding that quantumness allows us to outperform the classical world, which means that quantum objects are resources. The new technological developments, which are now becoming commercially available, demand a rigorous and thorough understanding of quantum resources, in order to quantify them precisely, and optimize their use. This is accomplished with the framework of resource theories [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. Resource theories are based on the simple observation that agents in a laboratory can only perform a subset of all possible deterministic quantum operations due to the various constraints on the agents and the systems involved. For instance, when two agents, Alice and Bob, are physically separated, they may not have a reliable means to exchange quantum systems, limiting them to local operations and classical communication (LOCC) [5, 6, 7]. Such allowed operations are called free.
@@
Remarkably, one notices that in each of these instances of resource theories a channel is considered free if and only if its renormalized Choi matrix is a free state. We call resource theories that exhibit this property Choi-defined resource theories. This is a highly valuable property: Every problem associated with free quantum channels can be converted into a problem involving free states.
Given the significance of Choi-defined resource theories and their computational convenience one wonders when it is possible to construct such resource theories from a set of free states. In this letter, we answer this question by providing necessary and sufficient conditions on the set of free states. Moreover, we show that when such a construction is possible, the resulting set of free operations coincides with the set of CRNG operations associated with those free states. These results provide a constructive definition of CRNG operations and generalize what has been already observed in the resource theories of magic, imaginarity, NPT, SEP, and non-negativity of quantum amplitudes. Next, we introduce resource measures, a complete family of monotones, and conversion distances in CDRTs, which can all be computed with conic linear programs (CLP) [36, 49] whenever the set of free states is convex and closed, or even with semidefinite programs (SDP) [50].
@@
    parallel composition of free channels is free.
@@
The renormalized Choi state is not the maximally entangled state in the resource theories of entanglement, SEP entanglement, or NPT entanglement. Indeed, in these resource theories, each system is a pair of physical systems [2], i.e., $S\coloneqq AB$, and $S^{\prime}\coloneqq A^{\prime}B^{\prime}$ and the spatial separation is between systems $AA^{\prime}$ and $BB^{\prime}$. Let {|$\Set{\Ket{a}_{A}}$ and {|$\Set{\Ket{b}_{B}}$ be orthogonal basis for $A$ and $B$, respectively, then
$\frac{1}{d_{S}}\Phi_{SS^{\prime}}=\frac{1}{d_{A}d_{B}}\sum_{a,b,\tilde{a},\tilde{b}}|ab\rangle\!\langle\tilde{a}\tilde{b}|_{AB}\otimes|ab\rangle\!\langle\tilde{a}\tilde{b}|_{A^{\prime}B^{\prime}}.$ (4)
@@
$\frac{1}{d_{A}d_{B}}\sum_{a,\tilde{a}}|aa\rangle\!\langle\tilde{a}\tilde{a}|_{AA^{\prime}}\otimes\sum_{b,\tilde{b}}|bb\rangle\!\langle\tilde{b}\tilde{b}|_{BB^{\prime}}$ (5)
$=\frac{1}{d_{A}}\Phi_{AA^{\prime}}\otimes\frac{1}{d_{B}}\Phi_{BB^{\prime}}.$
@@
§ V.  Examples of CDRTs
As mentioned in the introduction, known results about the resource theories of SEP, PPT, magic, imaginarity, and non-negative quantum amplitudes imply that these resource theories are CDRTs and we show in the SM that their sets of free states satisfy the conditions of Theorem 3. The attempt to construct a CDRT from $k$-extendible states made in Ref. [56] did not result in a proper resource theory because $k$-extendible states do not satisfy condition 2 of Theorem 3. The resource theory of $k$-unextendibility was later defined using different free operations [57].
Of more interest are those resource theories where the identification of free operations with CD operation has not been already established. For example, to the best of our knowledge, there is no explicit result yet in this direction for the resource theory of asymmetry [58, 14]. In the SM, we show that a resource theory of asymmetry is a CDRT if and only if the renormalized Choi state is free, which is equivalent to asking that all the unitary representations are real (cf. Supplementary Information of Ref. [59]).
§ VI.  Optimization problems in CDRTs
In CDRTs, every optimization problem over the set of free channels can be converted into an optimization problem over the set of free states. That is,
$\max_{\mathcal{M}_{A\to B}\in\mathfrak{F}(A\to B)}f(\mathcal{M}_{A\to B})$ (7)
becomes
$\max_{\begin{subarray}{c}\mu_{BA^{\prime}}\in\mathfrak{F}(BA^{\prime}),\\ d_{A^{\prime}}\operatorname{Tr}_{B}\mu_{BA^{\prime}}=\operatorname{\mathds{1}}_{A^{\prime}}\end{subarray}}\tilde{f}(\mu_{BA^{\prime}}).$ (8)
Here, $f$ is a function from the set of quantum channels to the $\mathbb{R}$, $\mathfrak{F}(A\to B)$ and $\mathfrak{F}(BA^{\prime})$ are the sets of free channels from $A$ to $B$ and free states on $BA^{\prime}$, and $\tilde{f}$ is the composition of $f$ with the inverse Choi isomorphism defined in Eq. (2), and it satisfies $\tilde{f}(\mu_{BA^{\prime}})=f(\mathcal{M}_{A\to B})$ whenever $\mu_{BA^{\prime}}$ is the renormalized Choi matrix of $\mathcal{M}_{A\to B}$. Moreover, if the set of free states is convex and $\tilde{f}$ is linear, the optimization problem can be expressed as a CLP (see SM for details) or even as an SDP.
If one has a set of free states that does not satisfy the conditions of Theorem 3, one can still consider the CRNG operations. However, in this case, there is no universal way of expressing the constraints in optimization problems as a single constraint in terms of free states.
§ VII.  Quantifying resources in CDRTs
@@
when $\operatorname{supp}\rho_{A}\subseteq\operatorname{supp}\sigma_{A}$ and +$\infty$+ ∞ otherwise. The resource measure based on the max-relative entropy is
$D_{\max}^{\mathfrak{F}}(\rho_{A})=\inf_{\omega_{A}\in\mathfrak{F}(A)}D_{\max}(\rho_{A}\|\omega_{A}).$ (10)
The extension of these quantities to quantum channels are [61, 62]
$D_{\max}(\mathcal{M}_{A\to B}\|\mathcal{N}_{A\to B})=$ (11)
$\quad\sup_{\rho_{RA}}D_{\max}(\mathcal{I}_{R}\otimes\mathcal{M}_{A\to B}(\rho_{RA})\|\mathcal{I}_{R}\otimes\mathcal{N}_{A\to B}(\rho_{RA})),$
$D_{\max}^{\mathfrak{F}}(\mathcal{M}_{A\to B})=\inf_{\mathcal{N}_{A\to B}\in\mathfrak{F}(A\to B)}D_{\max}(\mathcal{M}_{A\to B}\|\mathcal{N}_{A\to B}).$
Notably, the max-relative entropy for quantum channels satisfies [63]
$D_{\max}(\mathcal{M}_{A\to B}\|\mathcal{N}_{A\to B})=D_{\max}(\mu_{BA^{\prime}}\|\nu_{BA^{\prime}}),$ (12)
where $\mu_{BA^{\prime}}$ and $\nu_{BA^{\prime}}$ are the renormalized Choi matrices of $\mathcal{M}_{A\to B}$ and $\mathcal{N}_{A\to B}$, respectively. In a CDRT, this property is particularly relevant because it allows us to write the optimization problem for $D_{\max}^{\mathfrak{F}}(\mathcal{M}_{A\to B})$ as
$D_{\max}^{\mathfrak{F}}(\mathcal{M}_{A\to B})=\inf_{\begin{subarray}{c}\omega_{BA^{\prime}}\in\mathfrak{F}(BA^{\prime}),\\ d_{A^{\prime}}\operatorname{Tr}_{B}\omega_{BA^{\prime}}=\operatorname{\mathds{1}}_{A^{\prime}}\end{subarray}}D_{\max}(\mu_{BA^{\prime}}\|\omega_{BA^{\prime}}).$ (13)
Notice that, in every CDRT, $D_{\max}^{\mathfrak{F}}$ is finite for both states and channels because the set of free states always contain a full rank state, see Corollary 4. The same is not true for all resource theories.
As a direct consequence of the results in Ref. [63], the resource measures based on the max-relative entropy satisfy
@@
for all states $\rho_{A}$ and channels $\mathcal{M}_{A\to B}$. For the reasons mentioned above, this inequality, which is valid for all resource theories, is particularly meaningful in the context of CDRTs. When we interpret $D^{\mathfrak{F}}$ as the value of a resource, it implies that value of the output of a quantum operation applied to a quantum state is always less than the sum of the values of the operation and the input state. Equivalently, for every $\rho_{A}$, $\sigma_{B}$, and $\mathcal{M}_{A\to B}$ such that $\sigma_{B}=\mathcal{M}_{A\to B}(\rho)$, the cost of preparing $\sigma_{B}$ is always less or equal than the cost of preparing $\rho_{A}$ and then converting it to $\sigma_{B}$ with the channel $\mathcal{M}_{A\to B}$. An inequality similar to the one in Eq. (14) based on the quantum relative entropy and its regularization was derived in Ref. [64].
§ VIII.  Resource conversion in CDRTs
@@
Being a complete family of monotones, it satisfies $f_{\tau_{B}}(\rho_{A})\geq f_{\tau_{B}}(\sigma_{B})$ for all density matrices $\tau_{B}$ if and only if $\rho_{A}$ can be converted to $\sigma_{B}$ with a free operation.
@@
$=\inf_{\begin{subarray}{c}\mu_{BA^{\prime}}\in\mathfrak{F}(BA^{\prime}),\\ d_{A^{\prime}}\operatorname{Tr}_{B}\mu_{BA^{\prime}}=\operatorname{\mathds{1}}_{A^{\prime}}\end{subarray}}D(\sigma_{B},d_{A}\mu_{B\mathord{:}A^{\prime}}*\rho_{A}),$
where $D(\cdot,\cdot)$ is a function on quantum states that satisfies the data processing inequality [1].
§ IX.  Conclusions
In this work, we identified a key property of the resource theories of magic, imaginarity, SEP, NPT, and non-negativity of quantum amplitudes, which is that a channel is considered free if and only if its suitably renormalized Choi matrix is a free state. We presented a new way to construct a resource theory from the set of free states based on the Choi isomorphism and we derived which conditions on the free states are necessary and sufficient for constructing a mathematically consistent Choi-defined resource theory. Our interest is driven by the significant advantages that such a property guarantees. These advantages stem from the fact that every question in Choi-defined resource theories can be naturally framed as a question involving free states alone, which often have a well-understood structure. Indeed, we introduce several resource quantifiers which can all be computed with CLPs if the set of free states is convex and closed. Moreover, we showed that in any CDRT free operations are all CRNG operations. This provides a new constructive way to define CRNG operations when the set of free states satisfies the conditions of Theorem 3.
@@ Acknowledgements.
The authors are grateful to Matt Wilson for the several in-depth discussions on the subject of this letter and to Gilad Gour and Nathaniel Johnston for the interesting conversations about this research. E. Z. acknowledges support from the Eric Milner’s Graduate Scholarship, the Alberta Graduate Excellence Scholarship (AGES) and from the Eyes High International Doctoral Scholarship. C. M. S. acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) through the Discovery Grant “The power of quantum resources” RGPIN-2022-03025 and the Discovery Launch Supplement DGECR-2022-00119.